OpenStudy (haleyelizabeth2017):

I need an example of a rational expression that is equivalent with its simplified forms. I do not need anything else but an example I could use. :)

OpenStudy (triciaal):

example 3/5 example (3x + 2)/ 8 = 3x/8 + 1/4 example 24/3 = 8/1 = 8

OpenStudy (triciaal):

a rational expression means can be expressed as a ratio of 2 integers

OpenStudy (haleyelizabeth2017):

Okay thank you!

OpenStudy (triciaal):

welcome